Commercial paper. ​

Criss-Cross Manufacturers will issue commercial paper for a​ short-term cash inflow. The paper is for 91 days and has a face value of $50,000​, and the company anticipates it will sell at 96.8% of par value. ​ Criss-Cross wants to raise $3,500,000. What is the cost of this borrowing​ (annual terms)? How many​ "papers" will it​ sell?

What is the cost of this borrowing​ (three-month interest​ rate)?
Answer #1

How many​ "papers" will it​ sell?
=3500000/(50000*96.8%)
=72.314050

What is the cost of this borrowing​ (three-month interest​rate)?
=100/96.8-1
=3.30579%

What is the cost of this borrowing​ (annual terms)?
=3.30579%*4
=13.22316%

What is the effective cost of this borrowing​ (annual terms)?
=(1+3.30579%)^4-1
=13.89342%